A capacitor is made of two very large disks of rad
Answer #1

we know that

E = V/d = sigma /epsilon

80000 = sigma /epsilon

7.08x 10^-7 = sigma

sigma = Q/pi R^2

R = 1

Q = 2.224 x 10^-6 C ( answer)